The area of the rectangular playground enclosure at South school is 500 square meters. The length of the playground is 5 meters longer than the width. Find the dimensions of the playground in meters
Answer by stanbon(75887)

The area of the rectangular playground enclosure at South school is 500 square meters. The length of the playground is 5 meters longer than the width. Find the dimensions of the playground in meters
-----------------------------------
Area = L*W = 500
Length = W+5
----
Substitute and solve for "W":
(W+5)W = 500
w^2+5w-500 = 0
(w-20)(w+25) = 0
Positive solution:
Width = 20 meters
Length = W+5 = 25 meters
-----------------------------
Cheers,
Stan H.